Circular No. 5721

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 EA = 1984 AJ J. V. Scotti and D. L. Rabinowitz, Lunar and Planetary Laboratory,

report the following precise positions of an unusual asteroidal object discovered by Scotti with the 0.9-m Spacewatch telescope at Kitt Peak:

1993 UT  α_{2000}  δ_{2000}  V
Mar. 3.32749  12^{h}46^{m}18.^{s}07  +5^{o}05′29′′.6  19.3
3.33579  12 46 17.33  +5 05 36.0  19.3
3.34368  12 46 16.63  +5 05 42.3  19.2
3.43196  12 46 08.72  +5 06 51.8  19.1
3.44190  12 46 07.83  +5 06 59.7  19.3
3.44932  12 46 07.15  +5 07 05.7  19.2
4.23291  12 44 58.84  +5 17 26.3  19.2
4.23927  12 44 58.27  +5 17 31.3  20.1
4.24553  12 44 57.69  +5 17 36.3  19.5
19.25524  12 18 50.64  +8 44 41.4  19.0
19.26678  12 18 49.26  +8 44 50.6
19.27643  12 18 48.11  +8 44 58.6

Rabinowitz points out that his preliminary orbit determination indicates brightening to V = 16 during an approach to within 0.2 AU of the earth in Jan. 1994. The undersigned has identified the object with 1984 AJ, discovered by C. T. Kowal with the 1.2-m Oschin Schmidt at Palomar on 1984 Jan. 10, and the following orbit satisfies the above observations and the ends of the trail of the mag 15 image, measured by S. Swanson on Kowal’s 20-min exposure (cf. MPC 8510):

Epoch = 1993 Aug. 1.0 TT

T = 1993 Nov. 18.3350 TT  ω = 258.^{o}5931
e = 0.585640  Ω = 97.2466 }2000.0
q = 0.527056 AU  i = 5.0549
a = 1.271974 AU n^{o} = 0.6870471 P = 1.435 years
1993 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Mar. 24 12^{h}09.^{m}57 + 9^{o}47.^{′}8 1.019 2.006 168.^{o}8 5.^{o}5 19.0
Apr. 3 11 50.22 +11 42.8 1.033 1.994 157.6 11.0 19.3
13 11 33.12 +13 04.0 1.075 1.978 144.4 17.2 19.5
23 11 19.92 +13 47.7 1.138 1.957 131.8 22.5 19.8
May 3 11 11.36 +13 56.9 1.216 1.933 120.4 26.7 20.1

1993 March 21 Brian G. Marsden